# Optical absorption and luminescence of -LiV2O5 from the Bethe Salpeter Equation

### v1

@@ IV Discussion of the cathodoluminscence
Walker et al[4] studied changes in the electronic structure of LixV2O5 using depth resolved CL. Each of their peaks shows considerable fine structure that evolves with depth from bulk to surface and is also shown to be sensitive to remote oxygen plasma treatment, indicating that O-vacancies play a role. It is here not our intent to interpret these fine changes but rather to understand the broad main features and their changes between pure V2O5 and LiV2O5. Their spectrum for pure V2O5 shows a first peak just below 2 eV with fine structure at 1.8 and 1.9 eV, followed by a smaller peak near 3 eV and merging peaks at 3.6 and 4.1 eV. The main change upon lithiation is that the feature near 2 eV is strongly suppressed and essentially absent but can be partially restored by delithiation. These authors interpreted the feature near 2 eV as resulting from recombination of V- like split-off conduction electrons recombining with holes at the VBM while the higher peaks were associated with other V- orbitals, such as and . In a CL experiment electron hole pairs are first created by a high energy electron impact and then relax to the lowest states near the band edges before they radiatively recombine and emit light. So, first, it is not clear in this interpretation why the observed features are associated with density of states features in the conduction band only while ignoring the valence band structure and assuming that all holes have relaxed to the VBM. This would require one to assume a much faster hole than electron relaxation, for which there is no direct evidence, to the best of our knowledge. Second, the feature below 2 eV is about 0.5 below the aborption onset which was found to be near 2.35 eV by Kenny and Kannewurf[16]. Furthermore, adding Li would partially fill the split-off band and while one could imagine this to suppress absorption to these states, it would not really affect luminescent recombination between the split-off band and the VBM. The main observation of a decrease in the CL peak below 2 eV thus remains unexplained. Here we propose an alternative explanation.
@@
First, luminescence and absorption are proportional to but in luminescence the lower energy region of the spetrum should be emphasized by the carrier relaxation toward the lowest electron and hole states. Furthermore, spontaneous luminescence is proportional to . Second, the optical absorption in pure V2O5 has recently been shown to be dominated by very strongly bound excitons [17, 18]. The first strong absorption occurs at 3 eV followed by peaks at 4 eV and 5-6 eV with the quasiparticle gap at 3.8 eV. We thus associate the CL peaks between 3-4 eV with the direct excitons. The absorption onset of V2O5 near 2.3-2.6 eV on the other hand is still not entirely understood but is likely associated with indirect excitons and/or a dark exciton becoming activated by phonons or disorder as discussed in [38, 18]. But while there can be a Stokes shift between absorption and emission in localized state systems related to different structural relaxation in the ground and excited states, the CL feature at 1.8-1.9 eV is clearly not the same as the onset of absorption. We associate it with a polaronic emission or alternatively a self-trapped exciton. In fact, it has been shown by various calculations [20, 21, 22] that strong electron-phonon coupling leads to the formation of small electron polarons in V2O5. The added electrons are usually associated with O-vacancies, in particular the vanadyl-O-vacancy. Scanlon et al. [20] showed that O-vacancies produce defect states at about 1 eV above the VBM similar to Li interstitials when a DFT calculation is used which favors the formation of localized states and polaronic deformation. This was confirmed by Ngamwongwan et al. [21] and Watthaisong et al. [22] who showed that self-trapping of excitons ccurs near an added electron even without the presence of O-vacancies or nearby Li interstitials. The self-trapped polaron is found near the middle of the gap at about 1 eV above the VBM, while the O-vacancy related states depend on which type of O is removed. We thus propose that the CL feature at 1.8-1.9 eV is associated with recombination of a self-trapped electron with holes at the VBM induced in the CL experiment. We can also view this as a self-trapped indirect exciton. It is not straightforward to extract the energy position of this band from the GGA+U calculations of [20, 21, 22] as all of these miss the strong increase in quasiparticle gap we find in our QS calculations. However, we may assume that the polaron band lies close to the position of our Li induced band in LiV2O5 which is indeed about 2 eV above the VBM. While in our discussion above, this was associated with the spin splitting of the split-off band, the relaxation of the atoms near the Li which tilt the O-pyramids near each V toward the Li play also a decisive role in this feature being shifted down toward the middle of the gap. So, then why is this polaronic emission band reduced upon lithiation? At first one might think that the Li brings in more electrons and hence more polaron like states. However, there is now also a very strong absorption at about the same energy as the emission peak. We therefore propose that the emission is reduced by the strong absorption at this same energy. In other words, we propose that it it is the phenomenon of self-reabsorption that reduces the cathodolumnescence at 1.8 eV.
@@ V Conclusions
In this paper we showed that Li electron donation in -LiV2O5 leads to a splin splitting of the split-off conduction band which creates an occupied band in the middle of the gap. This then leads to strong optical absorption between equal spin bands with intra-atomic V- character. Essentially transitions occur between antisymmetric and symmetric linear combinations of the orbitals on the two V across a V-O-V rung, which form strongly localized molecular type states. This strong absorption is polarized exclusively along the direction. We propose a different interpretation of the CL experiments of Walker et al. [4] with the main peaks at 3-4 eV being associated with the charge transfer direct excitons of V2O5 and the 1.8 eV feature as an electron self-trapped polaron emission which is suppressed in LiV2O5 by self-absorption.

### v5

@@ IV Discussion of the cathodoluminscence
While our main results pertain to the absorption, we here wish to comment on the related cathodoluminescene (CL) spectra, which are the main experimental data on optical properties available at this time. Walker et al[4] studied changes in the electronic structure of LixV2O5 using depth resolved CL. Each of their peaks shows considerable fine structure that evolves with depth from bulk to surface and is also shown to be sensitive to remote oxygen plasma treatment, indicating that O-vacancies play a role. It is here not our intent to interpret these fine changes but rather to understand the broad main features and their changes between pure V2O5 and LiV2O5. Their spectrum for pure V2O5 , which is here reproduced approximately in Fig. 7 by digitizing the data from their figure, shows a first peak just below 2 eV with fine structure at 1.8 and 1.9 eV, followed by a smaller peak near 3 eV and merging peaks at 3.6 and 4.1 eV. The main change upon lithiation is that the feature near 2 eV is strongly suppressed and essentially absent but can be partially restored by delithiation. These authors interpreted the feature near 2 eV as resulting from recombination of V- like split-off conduction electrons recombining with holes at the VBM while the higher peaks were associated with other V- orbitals, such as and .
Considering first Walker et al. ’s interpretation, it is not clear why the observed features are associated with density of states features in the conduction band only while ignoring the valence band structure and assuming that all holes have relaxed to the VBM. This would require one to assume a much faster hole than electron relaxation, for which there is no direct evidence, to the best of our knowledge. Second, the feature below 2 eV is about 0.5 below the indirect absorption onset which was found to be near 2.35 eV by Kenny and Kannewurf[16]. It is thus clearly not a pure band to band transition or related band gap exciton but may be related to defects. Furthermore, adding Li would partially fill the split-off band and while one could imagine this to suppress absorption to these states, it would not really affect luminescent recombination between the split-off band and the VBM. The main observation of a decrease in the CL peak below 2 eV thus remains unexplained. Here we propose an alternative explanation.
Unfortunately, the interpretation of CL spectra is rather complex and a fully first-principles approach is not yet available. In a CL experiment electron hole pairs are first created by a high energy electron impact and then the electrons and holes relax to their lowest states near the band edges by multiple phonon emissions before they radiatively recombine and emit light. That last process is also in competition with non-radiative recombination. The spontaneous luminescence rate arising from the perfect crystal energy bands bands is proportional to the corresponding absorption coefficient with and . Here, is the real part of the index of refraction and is Boltzmann’s constant. However, as the excited electrons relax they increasingly populate lower energy states and hence for a defective crystal CL will more prominently show recombination from the defect levels in the gap to which the excited electrons trickle down.
@@
The optical absorption in pure V2O5 has recently been shown to be dominated by very strongly bound excitons [17, 18]. The first strong absorption occurs at 3 eV followed by peaks at 4 eV and 5-6 eV with the quasiparticle gap at eV in the QS method used in [18] and at slightly higher energies eV for the quasiparticle gap and 3.4 eV for the lowest exciton in QS as seen in Fig. 3. We thus associate the CL peaks between 3-4 eV with the direct excitons and band gap emission. The absorption onset of V2O5 near 2.3-2.6 eV on the other hand is still not entirely understood but is likely associated with indirect excitons and/or a dark exciton becoming activated by phonons or disorder as discussed in [39, 18]. The CL feature at 1.8-1.9 eV is clearly well below the onset of absorption. Thus, it must rather be related to recombination from a defect level in the gap. The most prevalent defect levels in V2O5 are related to oxygen vacancies and self-trapped polarons. In fact both correspond to electrons trapped on a single V. We therefore tentatively associate the lowest CL peak with a emission from a self-trapped polaron recombining with a hole in the VB, which could alternatively be described as a self-trapped exciton. In fact, it has been shown by various calculations [20, 21, 22] that strong electron-phonon coupling leads to the formation of small electron polarons in V2O5. The added electrons are usually associated with O-vacancies, in particular the vanadyl-O-vacancy. Scanlon et al. [20] showed that O-vacancies produce defect states at about 1 eV above the VBM similar to Li interstitials when a DFT calculation is used which favors the formation of localized states and polaronic deformation. This was confirmed by Ngamwongwan et al. [21] and Watthaisong et al. [22] who showed that self-trapping of excitons occurs near an added electron even without the presence of O-vacancies or nearby Li interstitials. In their calculations, the self-trapped polaron is found near the middle of the gap at about 1 eV above the VBM, while the O-vacancy related states depend on which type of O is removed. It is not straightforward to extract the energy position of this band from the GGA+U calculations of [20, 21, 22] as all of these miss the strong increase in quasiparticle gap we find in our QS calculations. However, we may assume that the polaron band lies close to the position of our Li induced occupied band in LiV2O5 since this also corresponds to adding an electron which localizes on a V- state, although it would in the present case be spread over two V. It is indeed about 2 eV above the VBM. Localization on a single V would pull it slightly further down. While in our discussion above, this was associated with the spin splitting of the split-off band, the relaxation of the atoms near the Li which tilt the O-pyramids near each V toward the Li play also a decisive role in this feature being shifted down toward the middle of the gap. In summary, we associate the eV CL band with a transition from this mid gap level to the VBM.
So, then why is this polaronic emission band reduced upon lithiation? At first one might think that the Li brings in more electrons and hence more polaron-like states. However, as we showed in the main results of the previous section there is now also a very strong absorption at about the same energy as the polaron emission peak. While in V2O5 even with some defects, the absorption from the defect levels to the conduction band is negligible compared to the emission resulting from recombination with the VBM holes, this is no longer the case in LiV2O5, where we found as main result a very strong optical absorption peak also at 1.8 eV. In fact, it is about a factor 10 stronger than the lowest exciton peaks in V2O5 and this is due to the more localized character of this optical transition between occupied and empty bands both dominated by the V- orbitals compared to the transition from a V- state to a more delocalized O- VB state involved in the emission process. We therefore propose that the emission is reduced by the strong absorption at this same energy. In other words, we propose that it is the phenomenon of “self-reabsorption” that reduces the cathodoluminescence at 1.8 eV. Interestingly, the self-trapped polaron level lies at about the same energy as the Li induced occupied band and this level lies almost exactly mid-gap so that absorption from this level to the CBM occurs at almost the same energy as recombination with a hole from the VBM.
@@ V Conclusions
In this paper we showed that Li electron donation in -LiV2O5 leads to a spin splitting of the split-off conduction band which creates an occupied band in the middle of the gap. This then leads to strong optical absorption between equal spin bands with intra-atomic V- character. Essentially transitions occur between antisymmetric and symmetric linear combinations of the orbitals on the two V across a V-O-V rung, which form strongly localized molecular type states. This strong absorption is polarized exclusively along the direction. We propose a different interpretation of the CL experiments of Walker et al. [4] with the main peaks at 3-4 eV being associated with the charge transfer direct excitons of V2O5 and the 1.8 eV feature as an electron self-trapped polaron emission which is suppressed in LiV2O5 by self-absorption.
Acknowledgements.
This work was supported by the U.S. Department of Energy Basic Energy Sciences (DOE-BES) under Grant No. DE-SC0008933. Calculations made use of the High-Performance Computing Resource in the Core Facility for Advanced Research Computing at Case Western Reserve University and the Ohio Supercomputer Center. Data Availability The data that support the findings of this article are openly available[40].References
- Sucharitakul et al. [2017] S. Sucharitakul, G. Ye, W. R. L. Lambrecht, C. Bhandari, A. Gross, R. He, H. Poelman, and X. P. A. Gao, V2O5: A 2D van der Waals Oxide with Strong In-Plane Electrical and Optical Anisotropy, ACS Applied Materials & Interfaces 9, 23949 (2017), pMID: 28677951.
- Whittingham [1976] M. S. Whittingham, The role of ternary phases in cathode reactions, Journal of The Electrochemical Society 123, 315 (1976).
- Zavalij and Whittingham [1999] P. Y. Zavalij and M. S. Whittingham, Structural chemistry of vanadium oxides with open frameworks, Acta Crystallographica Section B 55, 627 (1999).
- Walker et al. [2020] M. J. Walker, A. Jarry, N. Pronin, J. Ballard, G. W. Rubloff, and L. J. Brillson, Nanoscale depth and lithiation dependence of v2o5 band structure by cathodoluminescence spectroscopy, J. Mater. Chem. A 8, 11800 (2020).
- Jarry et al. [2020] A. Jarry, M. Walker, S. Theodoru, L. J. Brillson, and G. W. Rubloff, Elucidating Structural Transformations in LixV2O5 Electrochromic Thin Films by Multimodal Spectroscopies, Chemistry of Materials 32, 7226 (2020).
- Levy et al. [2023] Z. Levy, V. C. Ferrari, P. Rosas, M. J. Walker, K. Duddella, M. Haseman, D. Stewart, G. Rubloff, and L. J. Brillson, Lithium Spatial Distribution and Split-Off Electronic Bands at Nanoscale V2O5/LiPON Interfaces, ACS Applied Energy Materials 6, 4538 (2023).
- Warecki et al. [2024] Z. Warecki, V. C. Ferrari, D. A. Robinson, J. D. Sugar, J. Lee, A. V. Ievlev, N. S. Kim, D. M. Stewart, S. B. Lee, P. Albertus, G. Rubloff, and A. A. Talin, Simultaneous Solid Electrolyte Deposition and Cathode Lithiation for Thin Film Batteries and Lithium Iontronic Devices, ACS Energy Letters 9, 2065 (2024).
- Gillette et al. [2016] E. I. Gillette, N. Kim, G. W. Rubloff, and S. B. Lee, Interconnected mesoporous V2O5 electrode: impact on lithium ion insertion rate, Phys. Chem. Chem. Phys. 18, 30605 (2016).
- Chen et al. [2013] X. Chen, E. Pomerantseva, K. Gregorczyk, R. Ghodssi, and G. Rubloff, Cathodic ALD V2O5 thin films for high-rate electrochemical energy storage, RSC Adv. 3, 4294 (2013).
- Yue and Liang [2017] Y. Yue and H. Liang, Micro- and Nano-Structured Vanadium Pentoxide (V2O5) for Electrodes of Lithium-Ion Batteries, Advanced Energy Materials 7, 1602545 (2017).
- Rocquefelte et al. [2003] X. Rocquefelte, F. Boucher, P. Gressier, and G. Ouvrard, First-principle study of the intercalation process in the lixv2o5 system, Chemistry of Materials 15, 1812 (2003).
- Eyert and Höck [1998] V. Eyert and K.-H. Höck, Electronic structure of : Role of octahedral deformations, Phys. Rev. B 57, 12727 (1998).
- De Jesus et al. [2016] L. R. De Jesus, G. A. Horrocks, Y. Liang, A. Parija, C. Jaye, L. Wangoh, J. Wang, D. A. Fischer, L. F. J. Piper, D. Prendergast, and S. Banerjee, Mapping polaronic states and lithiation gradients in individual V2O5 nanowires, Nature Communications 7, 12022 (2016).
- Horrocks et al. [2016] G. A. Horrocks, E. J. Braham, Y. Liang, L. R. De Jesus, J. Jude, J. M. Velázquez, D. Prendergast, and S. Banerjee, Vanadium K-Edge X-ray Absorption Spectroscopy as a Probe of the Heterogeneous Lithiation of V2O5: First-Principles Modeling and Principal Component Analysis, The Journal of Physical Chemistry C 120, 23922 (2016).
- Bhandari et al. [2015] C. Bhandari, W. R. L. Lambrecht, and M. van Schilfgaarde, Quasiparticle self-consistent calculations of the electronic band structure of bulk and monolayer , Phys. Rev. B 91, 125116 (2015).
- Kenny et al. [1966] N. Kenny, C. Kannewurf, and D. Whitmore, Optical absorption coefficients of vanadium pentoxide single crystals, Journal of Physics and Chemistry of Solids 27, 1237 (1966).
- Gorelov et al. [2022] V. Gorelov, L. Reining, M. Feneberg, R. Goldhahn, A. Schleife, W. R. L. Lambrecht, and M. Gatti, Delocalization of dark and bright excitons in flat-band materials and the optical properties of V2O5, npj Computational Materials 8, 94 (2022).
- Garcia et al. [2024] C. Garcia, S. K. Radha, S. Acharya, and W. R. L. Lambrecht, Quasiparticle band structure and excitonic optical response in bulk and monolayer, Phys. Rev. B 110, 085102 (2024).
- Ioffe and Patrina [1970] V. A. Ioffe and I. B. Patrina, Comparison of the Small-Polaron Theory with the Experimental Data of Current Transport in V2O5, Physica Status Solidi (b) 40, 389 (1970).
- Scanlon et al. [2008] D. O. Scanlon, A. Walsh, B. J. Morgan, and G. W. Watson, An ab initio Study of Reduction of V2O5 through the Formation of Oxygen Vacancies and Li Intercalation, The Journal of Physical Chemistry C 112, 9903 (2008).
- Ngamwongwan et al. [2021] L. Ngamwongwan, I. Fongkaew, S. Jungthawan, P. Hirunsit, S. Limpijumnong, and S. Suthirakun, Electronic and thermodynamic properties of native point defects in V2O5: a first-principles study, Phys. Chem. Chem. Phys. 23, 11374 (2021).
- Watthaisong et al. [2019] P. Watthaisong, S. Jungthawan, P. Hirunsit, and S. Suthirakun, Transport properties of electron small polarons in a V2O5 cathode of Li-ion batteries: a computational study, RSC Adv. 9, 19483 (2019).
- Kotani et al. [2007] T. Kotani, M. van Schilfgaarde, and S. V. Faleev, Quasiparticle self-consistent GW method: A basis for the independent-particle approximation, Phys.Rev. B 76, 165106 (2007).
- Cunningham et al. [2018] B. Cunningham, M. Grüning, P. Azarhoosh, D. Pashov, and M. van Schilfgaarde, Effect of ladder diagrams on optical absorption spectra in a quasiparticle self-consistent framework, Phys. Rev. Materials 2, 034603 (2018).
- Cunningham et al. [2023] B. Cunningham, M. Grüning, D. Pashov, and M. van Schilfgaarde, : Quasiparticle self-consistent with ladder diagrams in , Phys. Rev. B 108, 165104 (2023).
- Pashov et al. [2019] D. Pashov, S. Acharya, W. R. Lambrecht, J. Jackson, K. D. Belashchenko, A. Chantis, F. Jamet, and M. van Schilfgaarde, Questaal: A package of electronic structure methods based on the linear muffin-tin orbital technique, Computer Physics Communications , 107065 (2019).
- Bhandari and Lambrecht [2015] C. Bhandari and W. R. L. Lambrecht, Electronic and magnetic properties of electron-doped and , Phys. Rev. B 92, 125133 (2015).
- Delmas et al. [1994] C. Delmas, H. Cognac-Auradou, J. Cocciantelli, M. Ménétrier, and J. Doumerc, The LixV2O5 system: An overview of the structure modifications induced by the lithium intercalation, Solid State Ionics 69, 257 (1994).
- Galy [1992] J. Galy, Vanadium pentoxide and vanadium oxide bronzes—Structural chemistry of single (S) and double (D) layer MxV2O5 phases, Journal of Solid State Chemistry 100, 229 (1992).
- van Schilfgaarde et al. [2006] M. van Schilfgaarde, T. Kotani, and S. Faleev, Quasiparticle Self-Consistent Theory, Phys. Rev. Lett. 96, 226402 (2006).
- Hedin [1965] L. Hedin, New method for calculating the one-particle green’s function with application to the electron-gas problem, Phys. Rev. 139, A796 (1965).
- Hedin and Lundqvist [1969] L. Hedin and S. Lundqvist, Effects of electron-electron and electron-phonon interactions on the one-electron states of solids, in Solid State Physics, Advanced in Research and Applications, Vol. 23, edited by F. Seitz, D. Turnbull, and H. Ehrenreich (Academic Press, New York, 1969) pp. 1–181.
- Ismail-Beigi [2017] S. Ismail-Beigi, Justifying quasiparticle self-consistent schemes via gradient optimization in baym–kadanoff theory, Journal of Physics: Condensed Matter 29, 385501 (2017).
- Aryasetiawan and Gunnarsson [1994] F. Aryasetiawan and O. Gunnarsson, Product-basis method for calculating dielectric matrices, Phys. Rev. B 49, 16214 (1994).
- Onida et al. [2002] G. Onida, L. Reining, and A. Rubio, Electronic excitations: density-functional versus many-body Green’s-function approaches, Rev. Mod. Phys. 74, 601 (2002).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized Gradient Approximation Made Simple, Phys. Rev. Lett. 77, 3865 (1996).
- Atzkern et al. [2001] S. Atzkern, M. Knupfer, M. S. Golden, J. Fink, A. N. Yaresko, V. N. Antonov, A. Hübsch, C. Waidacher, K. W. Becker, W. von der Linden, G. Obermeier, and S. Horn, One-dimensional dynamics of the d electrons in , Phys. Rev. B 63, 165113 (2001).
- Bhandari and Lambrecht [2014] C. Bhandari and W. R. L. Lambrecht, Phonons and related spectra in bulk and monolayer , Phys. Rev. B 89, 045109 (2014).
- Gorelov et al. [2023] V. Gorelov, L. Reining, W. R. L. Lambrecht, and M. Gatti, Robustness of electronic screening effects in electron spectroscopies: Example of , Phys. Rev. B 107, 075101 (2023).
- [40] https://github.com/Electronic-Structure-Group/LiV2O5_Optics.